A satellite orbits Earth 1,250,000 m above the Earth’s surface. What is the satellite’s orbital speed? 


1
Expert's answer
2021-04-30T11:24:29-0400
"v=\sqrt{\frac{GM}{r}}\\v=\sqrt{\frac{(6.67\cdot10^{-11})(5.97\cdot10^{24})}{1250000+6.59\cdot10^{6}}}=7130\frac{m}{s}"